OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form.

OpenStudy (texaschic101):

first find the slope....do you know how to do that ?

OpenStudy (anonymous):

I'm not sure if I'm right but 3/1 for the slope

OpenStudy (texaschic101):

let me check.....hold on

OpenStudy (johnweldon1993):

You are correct for the slope

OpenStudy (texaschic101):

m = (y2 - y1) / (x2 - x1) m = (-3 - 6) / (0 - 3) m = -9/-3 m = 3....slope is 3

OpenStudy (texaschic101):

now you need to use y = mx + b m = 3 for x and y, you can use either of your points....I will use (3,6) y = mx + b 6 = 3(3) + b now solve for b.....b is your y intercept

OpenStudy (anonymous):

Thank You :)

OpenStudy (texaschic101):

what equation did you get ?